OpenStudy (anonymous):

How you solve 1/2x + 13 = 9

OpenStudy (anonymous):

1/2x=-4=>2x=0.25=>x=0.125

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }*x +13-9 = 0\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } x + 4 = 0\]

OpenStudy (anonymous):

How you get 0

OpenStudy (anonymous):

\[\frac{ 1 }{2 }x = -4\]

OpenStudy (anonymous):

anything you do to one side of an equation you must do to the other, in this case ubtract 9 from both sides

OpenStudy (anonymous):

Alright thanks

OpenStudy (anonymous):

3/4x - 9 = 27

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

1/2 X + 13= 9 X= 9-13 / 1/2 X =-4 / 1/2 X= -2